HI
THIS IS MY QUESTION THAT I NEED SOLVING
PETER AND LYN ARE CARRYING OUT AN EXPERIMENT TO ESTIMATE THE NUMBER OF DIFFERENT COLOURED BALLS IN A BAG.
*THEY TAKE IT IN TURNS TO DRAW A BALL FROM THE BAG
*THEY NOTE ITS COLOUR AND RECORD IT IN A TABLE
*THEY REPEAT THIS 10 TIMES
*THE BAG CONTAINS 20 BALLS IN TOTAL
HERE ARE THE RESULTS
BLUE 3
GREEN 2
RED 4
YELLOW 1
USE THE RESULTS TO ESTIMATE THE PROBABILITY OF DRAWING EACH COLOUR BALL FROM THE BAG?

Answer by stanbon(48568)

PETER AND LYN ARE CARRYING OUT AN EXPERIMENT TO ESTIMATE THE NUMBER OF DIFFERENT COLOURED BALLS IN A BAG.
*THEY TAKE IT IN TURNS TO DRAW A BALL FROM THE BAG
*THEY NOTE ITS COLOUR AND RECORD IT IN A TABLE
*THEY REPEAT THIS 10 TIMES
*THE BAG CONTAINS 20 BALLS IN TOTAL
HERE ARE THE RESULTS
BLUE 3
GREEN 2
RED 4
YELLOW 1
USE THE RESULTS TO ESTIMATE THE PROBABILITY OF DRAWING EACH COLOUR BALL FROM THE BAG?
-----------------
P(blue) = 3/10
P(Green) = 2/10
P(red)=4/10
P(yellow)=1/10
==================
cheers,
Stan H.